# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v2

@@ 3.3.3. Workers
3.4. Using the design space
All categories and subcategories of the design space, which emerged through open coding, exist in both the crowdsourcing workflow and LLM chaining literatures. This section covers how designers can use the design space’s interdependent tiers to build LLM chains with an objectives-centered approach. We also discuss gaps in the literature, suggesting future work.
To exemplify this process, we follow the design of two hypothetical workflows as running examples: Workflow A performing an accuracy-based task with high resource constraints, and Workflow C performing a creativity-based task with few resource constraints. By following an objectives based approach, the workflows traverse the design space differently.
3.4.1. Objectives and the quality-cost tradeoff.
With a task in mind, designers can define their outcome qualities and resource constraints. Defining these objectives also defines how the chain must respect a quality-cost tradeoff. Mechanisms for improving output quality can be more expensive (Yao et al., 2023; Huffaker et al., 2020) and take longer (Chang et al., 2017; Gouravajhala et al., 2018). Some workflows have parameters for tuning this quality-cost tradeoff, such as giving the user control over the number of redundant generations (Yao et al., 2023; Huffaker et al., 2020; Drapeau et al., 2016; Parameswaran et al., 2023), parallelization of work on different parts of the input (Luther et al., 2015; Zharikova et al., 2023; LaToza et al., 2014), beam search (Yao et al., 2023; Xie et al., 2023; Creswell and Shanahan, 2022) and decision-theoretic choices of the number of responses to obtain (Bragg et al., 2013; Alshaibani and Quinn, 2021; Dai et al., 2010). However, this tradeoff is inherent to building a workflow due to the cost of querying for responses, and it impacts the degree to which strategies can be implemented.
Workflow A must be more mindful of the quality-cost tradeoff and prioritize consistently correct outcome qualities, while Workflow C prioritizes unconstrained idea generation as its primary creativity-focused outcome quality.
3.4.2. Objective-informed strategy choices.
The objectives inform how to utilize strategies within given resource constraints. An accuracy-focused outcome quality will often use aggregated diverse responses (Liang et al., 2023; Wu et al., 2023a; Du et al., 2023) and focus quality control on avoiding cascading errors and factual inaccuracies (Little et al., 2010b, a; Xie et al., 2023). Workflows for high-stakes accuracy tasks, which prioritize consistency, transparency, and debuggability, can prefer rigid and task-specific designs (Reppert et al., 2023; Wu et al., 2022a, 2023a). With limited resources, prioritizing the quality control strategy over the diverse responses and workflow flexibility strategies better suits the needs of tasks requiring high accuracy.
Workflows for creative outcomes use response diversity to improve divergent thinking and enable richer user interactions (Bernstein et al., 2010), and they use quality control methods to prioritize the coherency and qualities of the output. For example, verification steps in creative tasks focus on whether or not outputs meet given quality requirements, rather than if they are correct (Huang et al., 2017; Christoforaki and Ipeirotis, 2014; Chilton et al., 2019; Huang et al., 2015). Flexible workflows provide the ability to run creative tasks without pre-defining all edge cases, although the work must be easily decomposed (Kittur et al., 2011). With limited resources, prioritizing reducing architecture rigidity, if possible, and sourcing a creative diversity of responses can support creative outcome quality goals.
When considering strategies informed by objectives, Workflow A prioritizes many verification steps for quality control, then ensures enough workflow rigidity for high consistency. With remaining resources, it can source diverse responses for accuracy. In contrast, Workflow C generates diverse responses as its primary strategy.
3.4.3. Connecting strategies to tactics.
With strategies in mind, designers then turn to tactics: how to implement and parameterize a workflow. Many connections between strategies and tactics are intuitive: redundant architectures source diverse responses, evaluate subtasks can be a verification step, and dynamic architectures create non-rigid workflows. Others are less intuitive. For example, consistency between parallel workflows serve as a verification step (Liem et al., 2011; Parameswaran et al., 2023), redundant architectures help with user revision (Bernstein et al., 2010), and communicative architectures source and evaluate diverse creative responses (Chilton et al., 2014; André et al., 2013; Retelny et al., 2014). Approaching strategies with objectives in mind also informs tactics. For example, sourcing diverse responses for a creative task with coherency as a quality goal would prioritize iterative, rather than parallel, redundant architectures.
Workflow A’s architecture uses rigid subtasks to ensure consistency, but, to increase efficiency, incorporates a dynamic element controlling the number of redundant generations based on available resources. Workflow A uses evaluate subtasks to verify correctness. Workflow C also uses redundant architectures to source diverse responses but instead resolves disagreements in those generations with transform subtasks to iteratively refine and improve the ideas.
3.4.4. Gaps between literatures in the design space.
Another utility of the design space is to compare literatures to see if there are elements of crowdsourcing that can better inform LLM chains. Some parts of the design space include more varied approaches in the crowdsourcing literature, including tasks’ outcome qualities, approaches to reducing workflow rigidity such as free-form communicative architectures, and subsets of approaches to user revision such as creating controllable outputs. Crowdsourcing research also studies a wider array of specific tactical designs, such as the effects of parallel and iterative redundant architectures with different parameters (André et al., 2014; Little et al., 2010a; Kittur et al., 2011; Goto et al., 2016; Luther et al., 2015; Huffaker et al., 2020). To take advantage of this extensive crowdsourcing workflow history, it would be useful to know if designers can incorporate findings from these studies into chaining.
Since LLMs are different from crowdworkers, it is unclear from the literature alone what will effectively transfer to chaining and what will not. Some studies begin to compare LLM chains and crowdsourcing workflows (Wu et al., 2023c; Parameswaran et al., 2023), but much remains unknown. To better understand the extent to which crowdsourcing can inform a more expansive and detailed design space for chaining, we adapt crowdsourcing workflows into LLM chains (Section 4). We then synthesize our findings from the literature and our case study experiences to propose what does (Section 5.1) and does not (Section 5.2) transfer from crowdsourcing to chaining and the resulting implications for future work.
@@ 4. Case studies
In order to better assess how crowdsourcing methods transfer to LLM chain development, we adapt crowdsourcing workflows to LLM chains in three case studies. We document insights from this process and map our decisions into the design space. The case studies foray into an underexplored part of the LLM space: supporting quality control through user revision in which the user can directly manipulate workflow outputs. With directly manipulable outputs, users can make “rapid, incremental, reversible operations whose impact on the object of interest is immediately visible” (Shneiderman, 1982), increasing control over output elements and helping users explore alternative results (Hutchins et al., 1985). By adapting this subset of the user revision strategy from crowdsourcing, we explore how and if such a concept can transfer to chains and what modifications are necessary. Prior non-chaining LLM work explores controllability as a helpful attribute (MacNeil et al., 2023; Kim et al., 2023a), so here we seek to incorporate it within chains.
We choose three popular crowdsourcing tasks that span the verifiable creativity and creativity sections of the design space quality objectives. We implement Cascade (Chilton et al., 2013) to build taxonomies with manipulable structures, Soylent (Bernstein et al., 2010) to shorten text to a variety of target lengths, and Mechanical Novel (Kim et al., 2017) to control the elements of a story. Diagrams of the workflows, as well as our adaptations to use LLMs, can be found in Figures 6, 8, and 10.
For each case study, we first document how the workflow was changed to incorporate LLMs and map those changes to the design space. Common limitations we faced include LLM context lengths, inconsistent output formatting, lack of diversity in model responses, and models struggling to perform the workflow-defined subtasks. We adapted to these limitations with more granular subtasks, more diverse prompts, comparative tasks, and hard-coded verification steps. Across all three case studies, we use different prompts at each parallel generation step to increase response diversity, update worker instructions through trial and error to get the right formatting from the LLM outputs, and structure voting evaluation subtasks to choose among numbered options to minimize the impact of hallucinations.
@@ 4.1. Case study 1: Taxonomy creation
For our first case study, we explore taxonomy creation, the task of organizing a set of items into a category tree to aid sensemaking. This task falls under the verifiable creativity aspect of the design space: categories need to be correct, but most items can fit into multiple categorizations of different levels of precision. In our design, we focus on the ability to control the level of granularity of these categorizations by controlling the taxonomy shape.
@@ 4.1.1. Workflow design
We adapt the Cascade (Chilton et al., 2013) crowdsourcing workflow (Figure 6), which takes as input a list of items and outputs a tree mapping items to categories. The workflow begins by selecting a subset of 32 items, generating category suggestions for each item, selecting the best categories among the generated options, and confirming which among all categories works for each item. The next step turns the item-category assignments into a taxonomy, deleting small categories and merging similar categories. The workflow then updates the categorizations for the entire list of items with this new assignment. The workflow iterates with a new subset of items until all items are categorized, building a progressively larger set of item categorizations.
To build Cascade-LLM, our version of Cascade powered by LLMs, we made multiple changes to the original workflow. We highlight some of the most salient modifications. To achieve diverse categorizations, we improved response diversity by giving different prompts to parallel generations and ensuring that every item is input into a category generation step. At first, the categories were too broad. Specifying category qualities via targeted prompts proved unsuccessful, so instead we added a programmatic validation step to delete categories that included “and,” “or,” commas, or slashes. As this hard-coded validation step worked, it was not necessary to include another LLM call. The original workflow gives workers sets of items for which to generate categories. However, the LLM would try categorizing multiple items at once, consistent with prior findings that LLMs struggle to disambiguate multiple subtasks in a prompt (Wu et al., 2023c; Madaan et al., 2023b). To increase diversity, we found the appropriate subtask design includes one item categorization. Finally, the original workflow also requires parameters for the minimum category size and a merge threshold. To support user revision through direct manipulation, we dynamically apply these parameters at the last taxonomization stage.
@@ 4.1.2. Evaluation setup
Datasets. Our experiments use the labels from the scenes in MIT Indoor Scenes (Quattoni and Torralba, 2009), object categories in MSCOCO (Lin et al., 2014b), and in CIFAR100 (Krizhevsky et al., 2009) datasets. These label sets are of size 67, 80, and 100, respectively. The original Cascade paper explores three datasets sourced from Quora.com question responses. Our LLM workflow is capable of processing these larger datasets, as only individual items need to be within the model’s context size; however, the Quora.com data exceeds the context limit for the zero-shot baselines.
@@ 4.1.5. Workflows have higher costs
Figure shows two flow charts, one for crowdsourcing Soylent and one for LLM chain Soylent. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: 1) the initial text splits into smaller chunks, 2) the Find and Fix steps are split into parallel edit, merge, and delete steps, and 3) automated format verification is added after the fix step.
@@ 4.2. Case study 2: Text shortening
The next case study is the task of shortening text. Unlike summarization, shortening the text involves making edits to create a condensed version of the text that attempts to preserve the full meaning. Shortening is a task with verifiable creativity, as shortened versions must stay faithful to the original text, but there are many options of what phrases to shorten and how to shorten them. By implementing a workflow to generate many shortening options, an end user can control the final length of a paragraph dynamically.
@@ 4.2.1. Workflow design
To shorten text we adapt the Find-Fix-Verify workflow (Figure 8) introduced in Soylent (Bernstein et al., 2010). In the “Find” step, crowdworkers select areas that might be shortened. If at least two workers agree on a text span, that span progresses to the “Fix” step. In the “Fix” step, crowdworkers make edits to shorten the text. In the final “Verify” step, more crowdworkers verify the possible changes by rejecting by majority vote any edits deemed grammatically incorrect or that change the meaning of the sentence. The workflow outputs edit options for multiple spans in the paragraph, which in turn may have multiple possible edits. Soylent defaults to showing the shortest possible option, but also provides a slider for the user to adjust the final output as desired.
We made multiple adjustments to produce our Soylent-LLM workflow. First, we ran into context limits and parallelization constraints with longer texts. In response, we apply a dynamic approach that splits the workflow into chunks of 10 sentences and applies the workflow to each. Second, Soylent allows workers to fix inputs by editing, merging, and deleting phrases. LLMs performed better when we split these tasks into separate prompts, so we further decompose Find and Fix into finer grained editing, merging, and deleting steps to get an appropriate subtask size. Finally, in the Find step the LLM would sometimes “hallucinate” phrases to shorten that do not exist in the text or would change quotes. To mitigate these errors, we add several hard-coded verification steps: we ensure that the phrase to replace exists in the input text, that the replacement is shorter, and that no quotes were altered or added.
@@ 4.2.2. Evaluation setup
Datasets. We use the XSUM summarization dataset (Narayan et al., 2018). We only use the input texts because our task is shortening, not matching gold-standard summarizations. We use 100 randomly sampled texts from XSUM, with the requirement that examples are more than 500 words long. For qualitative analysis, we use the texts from the Soylent evaluation (Bernstein et al., 2010). Through this dataset, we can compare the Soylent-LLM shortenings to Crowdsourced Soylent shortenings.
@@ 4.2.3. Workflows enable length manipulation
We first compare how well the various LLM approaches reach a desired output length. We represent target lengths through word counts of words. In initial experiments, we found that zero-shot baselines perform more competitively when asked to shorten to a given word count rather than to a length percentage and hence use word counts in targeted prompts. Nevertheless, we also measure accuracy in terms of a target percentage of length. We calculate the target percent length by dividing the target number of words by the input text’s number of words, and we calculate the actual percent length as the output text’s number of words divided by the input text’s number of words. We compare the percent difference (i.e., error) between the actual and target percent lengths.
We find that as text length increases, Soylent-LLM can more precisely meet the target length (Figure 9). When shortening text, the zero-shot prompt condition consistently outputs approximately word long shortenings (mean: , std: ). Thus, for longer target lengths, this method achieves lower performance. Zero-shot-target is similarly best at lengths close to words, but it still shows a lack of precision in reaching longer target lengths. Soylent-LLM and zero-shot-ffv both struggle to hit shorter targets. This result matches the Soylent paper, in which output texts were between 78-90% of the original text length (Bernstein et al., 2010). Soylent-LLM can more precisely hit the larger target lengths than any of the baselines.
We find similar trends measuring percent error with the percent length reduction required to meet the target length. Soylent-LLM and zero-shot-ffv can perform poorly with large reductions, in which the target % is a small fraction of the input text, while zero-shot and zero-shot-target perform worse with smaller reductions. In summary, Soylent-LLM does a better job of tightening long texts but is less well-suited to extreme shortenings that cross over into summarization. In addition, Soylent-LLM is unique in providing options among all these different targets without having to requery the model. Moreover, at each target length, there are often multiple possible output texts.
@@ 4.2.4. Workflow outputs have abrupt transitions but lower perplexity
Soylent-LLM has the better result of a lower perplexity (mean: , std: ) than zero-shot (mean: , std: ), zero-shot-target (mean: , std: ), and zero-shot-ffv (mean: , std: ). This result surprised us, as zero-shot is fully generated by a large language model fine-tuned to provide fluent responses.

### v3

@@ 6. Case studies
In order to better assess how crowdsourcing methods transfer to LLM chain development, we adapt crowdsourcing workflows to LLM chains in three case studies. This process of adapting workflows supports two investigations. First, it allows us to compare the performance of chaining to both zero-shot prompting and crowdsourcing. Second, through the process of adapting these workflows from crowdsourcing to chaining, we can see how the workflow design changes to accommodate for LLMs. These changes, found through practice, directly compare what tactics support the same strategies and objectives when using crowdworkers or LLMs.
To ensure we are suitably focused, we explore one strategy in more depth: supporting quality control through user guidance in which the user can directly manipulate workflow outputs. We focus on this strategy as it is implemented in crowdsourcing work, but it is underexplored in the LLM chaining space. Prior non-chaining LLM work explores controllability as a helpful attribute (MacNeil et al., 2023; Kim et al., 2023a), so here we seek to incorporate it within chains. With directly manipulable outputs, users can make “rapid, incremental, reversible operations whose impact on the object of interest is immediately visible” (Shneiderman, 1982), increasing control over output elements and helping users explore alternative results (Hutchins et al., 1985). By adapting this subset of the user guidance strategy from crowdsourcing, we explore how and if such a concept can transfer to chains, what modifications are necessary, and its relative advantages compared to controlling the outputs through zero-shot prompting.
We choose three popular crowdsourcing tasks that incorporate different degrees of the accuracy and creativity design space objectives. We implement Cascade (Chilton et al., 2013) to build taxonomies with manipulable structures, Soylent (Bernstein et al., 2010) to shorten text to a variety of target lengths, and Mechanical Novel (Kim et al., 2017) to control the elements of a story. Diagrams of the workflows, as well as our adaptations to use LLMs, can be found in Figures 13, 16, and 18.
For each case study, we first document how the workflow was changed to incorporate LLMs and map those changes to the design space. Common limitations we faced include LLM context lengths, inconsistent output formatting, lack of diversity in model responses, and models struggling to perform the workflow-defined subtasks. We adapted to these limitations with more granular subtasks, more diverse prompts, comparative tasks, and hard-coded validation steps. Across all three case studies, we use different prompts111https://github.com/madeleinegrunde/designingchains at each parallel generation step to increase response diversity, update worker instructions and context through trial and error to get the right output formatting, and structure voting evaluation subtasks to choose among numbered options to minimize hallucinations.
@@
Four taxonomies. The first with default parameters (merge threshold 0.75, minimum category size 2), contains a hierarchy of 4 layers and 23 categories. The second that is more likely to merge (merge threshold 0.5, min category size 2), contains 4 layers and 18 categories. The third that is a larger minimum category size (merge threshold 0.75, min category size 5) contains 3 layers and 13 categories). The fourth that is both more likely to merge and a larger min category size (merge threshold 0.5, min category size 5), contains 3 layers and 9 categories).
@@ 6.1. Case study 1: Taxonomy creation
For our first case study, we explore taxonomy creation, the task of organizing a set of items into a category tree to aid sensemaking. This task involves many elements of the accuracy objective, with some degree of creativity: categories need to be correct, but most items can fit into multiple categorizations of different levels of precision. We focus on the ability to control the granularity of these categorizations by controlling the taxonomy shape (Figure 15).
@@ 6.1.1. Workflow design
We adapt the Cascade (Chilton et al., 2013) crowdsourcing workflow (Figure 13), which takes as input a list of items and outputs a tree mapping items to categories. The workflow begins by selecting a subset of 32 items, generating category suggestions for each item, evaluating by selecting the best categories among the generated options, and evaluating by confirming which among all categories works for each item. The next step turns the item-category assignments into a taxonomy, deleting small categories and merging similar categories. The workflow then updates the categorizations for the entire list of items with this new assignment. The workflow redundantly iterates with new subsets until all items are categorized, building a growing set of categorizations.
To build Cascade-LLM, our version of Cascade powered by LLMs, we made multiple changes to the original workflow. We highlight some of the most salient modifications. To achieve diverse categorizations, we improved response diversity by giving different prompts to parallel generations and ensuring that every item is input into a category generation step. At first, the categories were too broad. Specifying category qualities via targeted prompts proved unsuccessful, so instead we added a programmatic validation step to delete categories that included “and,” “or,” commas, or slashes. As this hard-coded validation step worked, it was not necessary to include another LLM call. The original workflow gives workers sets of 8 items for which to generate 8 categories. However, the LLM would try categorizing multiple items at once, consistent with prior findings that LLMs struggle to disambiguate multiple subtasks in a prompt (Wu et al., 2023c; Madaan et al., 2023b). To increase diversity, we aligned subtask design to fit model capabilities with one item categorization. Finally, the original workflow also requires parameters for the minimum category size and a merge threshold. To support user guidance through direct manipulation, we dynamically apply these parameters at the last taxonomization stage.
@@ 6.1.2. Evaluation setup
Datasets. Our experiments use the labels from the scenes in MIT Indoor Scenes (Quattoni and Torralba, 2009), from object categories in MSCOCO (Lin et al., 2014b), and from CIFAR100 (Krizhevsky et al., 2009). These label sets are of size 67, 80, and 100, respectively. The original Cascade paper explores three datasets sourced from Quora.com question responses. Our LLM workflow is capable of processing these larger datasets, as only individual items need to be within the model’s context size; however, the Quora.com data exceeds the context limit for the zero-shot baselines.
@@ 6.1.5. Workflows have higher costs
Figure shows two flow charts, one for crowdsourcing Soylent and one for LLM chain Soylent. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: 1) the initial text splits into smaller chunks, 2) the Find and Fix steps are split into parallel edit, merge, and delete steps, and 3) automated format validation is added after the fix step.
@@ 6.2. Case study 2: Text shortening
The next case study is the task of shortening text. Unlike summarization, shortening the text involves making edits to create a condensed version of the text that attempts to preserve the full meaning. Shortening is a task on the creative side, but requiring elements of accuracy. Shortened versions must stay faithful to the original text, and there are many options of what phrases to shorten and how to shorten them. By implementing a workflow to generate many shortening options, an end user can control the final length of a paragraph dynamically.
@@ 6.2.1. Workflow design
To shorten text we adapt the Find-Fix-Verify workflow (Figure 16) introduced in Soylent (Bernstein et al., 2010). In the “Find” step, crowdworkers select areas that might be shortened. If at least two workers agree on a text span, that span progresses to the “Fix” step. In the “Fix” step, crowdworkers make edits to shorten the text. In the final “Verify” step, more crowdworkers verify the possible changes by rejecting by majority vote any edits deemed grammatically incorrect or that change the meaning of the sentence. These steps respectively map to the focus, transform, and evaluate subtask types in our design space. The workflow outputs edit options for multiple spans in the paragraph, which in turn may have multiple possible edits. Soylent defaults to showing the shortest possible option, but it also provides a length slider for the user to adjust the final output as desired.
We made multiple adjustments to produce our Soylent-LLM workflow. First, we ran into context limits and parallelization constraints with longer texts. In response, we apply a dynamic approach that splits the workflow into chunks of 10 sentences and applies the workflow to each. Second, Soylent allows workers to fix inputs by editing, merging, and deleting phrases. LLMs performed better when we split these tasks into separate prompts, so we further decompose Find and Fix into finer grained editing, merging, and deleting steps to get an align subtask size to the model’s capabilities. Finally, in the Find step the LLM would sometimes “hallucinate” phrases to shorten that do not exist in the text or would change quotes. To mitigate these errors, we add several hard-coded validation steps: we ensure that the phrase to replace exists in the input text, that the replacement is shorter, and that no quotes were altered or added.
@@ 6.2.2. Evaluation setup
Dataset. We use the XSUM summarization dataset (Narayan et al., 2018). We only use the input texts because our task is shortening, not matching gold-standard summarizations. We use 100 randomly sampled texts from XSUM, with the requirement that examples are more than 500 words long. For qualitative analysis, we use the texts from the Soylent evaluation (Bernstein et al., 2010). Through this dataset, we can compare the Soylent-LLM shortenings to Crowdsourced Soylent shortenings.
@@ 6.2.3. Workflows enable length manipulation
We first compare how well the various LLM approaches reach a desired output length. We represent target lengths through word counts of 100-500 words. In initial experiments, we found that zero-shot baselines perform more competitively when asked to shorten to a given word count rather than to a length percentage and hence use word counts in targeted prompts. Nevertheless, we also measure accuracy in terms of a target percentage of length. We calculate the target percent length by dividing the target number of words by the input text’s number of words, and we calculate the actual percent length as the output text’s number of words divided by the input text’s number of words. We compare the percent difference (i.e., error) between the actual and target percent lengths.
We find that as text length increases, Soylent-LLM can more precisely meet the target length (Figure 17). When shortening text, the zero-shot prompt condition consistently outputs approximately 100 word long shortenings (mean: 126.14, std: 52.63). Thus, for longer target lengths, this method achieves lower performance. Zero-shot-target is similarly best at lengths close to 100 words, but it still shows a lack of precision in reaching longer target lengths. Soylent-LLM and zero-shot-ffv both struggle to hit shorter targets. This result matches the Soylent paper, in which output texts were between 78-90% of the original text length (Bernstein et al., 2010). Soylent-LLM can more precisely hit the larger target lengths than any of the baselines.
We find similar trends measuring percent error with the percent length reduction required to meet the target length. Soylent-LLM and zero-shot-ffv can perform poorly with large reductions, in which the target % is a small fraction of the input text, while zero-shot and zero-shot-target perform worse with smaller reductions. In summary, Soylent-LLM does a better job of tightening long texts but is less well-suited to extreme shortenings that cross over into summarization. In addition, Soylent-LLM is unique in providing options among all targets without having to requery. Moreover, at each target length, there are often multiple possible output texts.
@@ 6.2.4. Workflow outputs have abrupt transitions but lower perplexity
Soylent-LLM has the better result of a lower perplexity (mean: 11.69, std: 3.77) than zero-shot (mean: 15.37, std: 5.73), zero-shot-target (mean: 12.31, std: 3.91), and zero-shot-ffv (mean: 16.19, std: 5.68). This result surprised us, as zero-shot is fully generated by a large language model fine-tuned to provide fluent responses.
@@
Figure shows two flow charts, one for crowdsourcing Mechanical Novel and one for LLM chain Mechanical Novel. Both follow the workflow described in the main text. Several adaptations are highlighted in the LLM chain: 1) additional validation steps in the reflect part of the workflow, 2) an automated length validation step in the revise part of the workflow 3) extra iterations through the scenes of the text as well as with different seeded texts.